OpenStudy (anonymous):

Can someone help me solve this line AB passes through A(-3, 0) and B(-6, 5). What is the equation of the line that passes through the origin and is parallel to line AB ?

OpenStudy (anonymous):

5x – 3y = 0 -x + 3y = 0 -5x – 3y = 0 3x + 5y = 0 -3x + 5y = 0